3. A jar contains 7 red, 2 green, and 4 pink marbles. If you pick one without looking, what is
the probability that the marble you pick will not be green?
Answer #1

No of marbles :

Red = 7

Green = 2

Pink = 4

Toal marbles = 13

P(green) = 2/13

P(not green) = 1- 2/13 = 11/13

ANS : Probability that the marble will not be green = 11/13 = 0.846